OpenStudy (anonymous):

find the coordinates of the midpoint of HX. H(-6.3,5.2) , X(1.8, -1)

OpenStudy (anonymous):

use the formula : xm = (x1+x2)/2 ym = (y1+y2)/2

OpenStudy (anonymous):

i know i dont understand

OpenStudy (anonymous):

from data given, x1=-6.3 , x2=1.8 , y1=5.2 and y2=-1 just plug of them to formula above

OpenStudy (anonymous):

-2.25 and 2.1 ?

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

yes, u are right